# A semi-analytical solution for the lubrication force between two spheres approaching in viscoelastic fluids described by the Oldroyd-B model under small Deborah numbers

### v1

@@ Abstract
Viscoelastic fluids play a critical role in various engineering and biological applications, where their lubrication properties are strongly influenced by relaxation times ranging from microseconds to minutes. Although the lubrication mechanism for Newtonian fluids is well-established, its extension to viscoelastic materials - particularly under squeezing flow conditions - remains less explored. This study presents a semi-analytical solution for the lubrication force between two spheres approaching in a Boger fluid under small Deborah numbers. Unlike previous works that assumed a Newtonian velocity field, we derive the velocity profile directly from the mass-momentum conservation and Oldroyd-B constitutive equations using lubrication theory and order-of-magnitude analysis techniques. Under steady-state conditions, viscoelasticity induces a marginal increase in the surface-to-surface normal force as a result of the increased pressure required to overcome the original resistance from the first normal-stress difference. Transient analyses reveal that the normal lubrication force is bounded by two Newtonian plateaus and is nonsymmetric as the spheres approach or separate. Our findings highlight the role of viscoelasticity in improving load capacity and provide new insight for modelling dense particle suspensions in Boger fluids, where short-range interactions dominate.
@@ 1 Introduction
Viscoelastic fluids are fundamental in various technological and biological contexts involving close-contact flows, such as lubrication of sliding surfaces, synthetic transport in living organisms, and synovial joint dynamics (Ahmed and Biancofiore, 2021; Urry, 1993; Ateshian, 2009). Their defining feature—a wide range of relaxation times—leads to complex nonlinear responses governed by normal stress differences (Zhang and Li, 2005; Nam et al., 2015). Despite extensive studies on Newtonian suspensions (Hoffman, 1972; Zarraga et al., 2000; Scirocco et al., 2020; Guazzelli and Pouliquen, 2018), suspensions in viscoelastic media remain less explored. Experiments with viscoelastic fluids, including Boger fluids characterized by a nearly constant shear viscosity and pronounced elasticity (Oldroyd and Wilson, 1950; Boger, 1977), reveal distinctive behaviours, such as early shear thickening, complex stress scaling, and matrix–microstructure interactions (Zarraga et al., 2001; Scirocco et al., 2004; Dai and Tanner, 2020). Suspensions in viscoelastic fluids also exhibit particle migration, chaining, and flow-induced structuring (Michele et al., 1997; Pasquino et al., 2013; Del Giudice et al., 2013). Microscopy studies in shear-thinning media show string-like particle alignment at low Weissenberg numbers, governed by fluid rheology and particle size (Pasquino et al., 2010). Settling experiments indicate that elasticity and shear-thinning enhance migration, unlike Newtonian fluids where such effects are suppressed (Scirocco et al., 2020; D’Avino et al., 2012). In confined microfluidic environments, particles can self-organise into ordered structures due to coupled inertial and viscoelastic forces, underscoring the significance of confinement and non-Newtonian flow features in particle-laden systems (Villone et al., 2011; Giudice and D’Avino, 2024).
@@
Numerical studies on viscoelastic suspensions have largely focused on dilute to semi-dilute regimes (volume fractions up to $\phi=0.3$), offering valuable insights into particle-laden viscoelastic flows through a variety of computational frameworks. Vázquez-Quesada et al. (2019) employed Smoothed Particle Hydrodynamics (SPH), incorporating a GENERIC-based viscoelastic model, to simulate rigid particle suspensions. They observed a plateau in relative viscosity at low Deborah numbers (De), followed by shear thickening at higher De due to extensional effects in the matrix. Jain and Shaqfeh (2021) used body-fitted (BF) and immersed-boundary (IB) methods to study Boger fluid suspensions, finding increases in per-particle viscosity and the first normal stress coefficient with strain, particle volume fraction, and Weissenberg number. Interestingly, fluid stress remained independent of $\phi$, while the stresslet rose due to enhanced hydrodynamic interactions. Villone et al. (2021) investigated immersed-boundary-method numerical simulations of small-amplitude oscillatory shear in shear-thinning and viscoelastic matrices, showing inertia-induced elasticity and Carreau number–dependent shear-thinning effects. Viscoelastic moduli varied with Deborah number and viscosity ratio, but not with the Giesekus mobility factor. Several methodologies have expanded the modelling tool set: Kuron et al. (2021) proposed a lattice Boltzmann method (LBM) for Oldroyd-B fluids, preserving stress and energy in complex flows. Matsuoka et al. (2021) and Yamamoto et al. (2021) applied the smoothed profile method (SPM), effectively predicting shear thickening and accommodating complex particulate systems. Zhang and Shaqfeh (2023) combined 3D finite volume simulations with experiments to study shear-thinning suspensions, linking reduced viscosity to diminished stresslet contributions and particle-induced stresses. Recent extensions of these frameworks include the lattice Boltzmann–smoothed profile hybrid method (LBM-SPM) by Lee (2023), and the thermally augmented SPM by Nakayama et al. (2024), which enables accurate micro-rheological analysis in thermally fluctuating viscoelastic fluids. Previous simulations of passive microrheology for a colloidal particle immersed in a thermodynamic consistent fluctuating Oldroyd-B fluid were already presented earlier in Vázquez-Quesada et al. (2012).
@@
Despite the significant numerical work performed in the dilute and semi-dilute regimes, examining dense suspensions where $\phi>0.3$ has been much less investigated. For Newtonian suspensions, recent advances in simulations of the rheology of concentrated suspensions—particularly dense, non-colloidal systems—highlight the crucial role of microstructure, contact mechanics, and interparticle forces in governing non-Newtonian flow phenomena (see Morris, 2009; Seto et al., 2013; Wyart and Cates, 2014; Tanner, 2018). To simulate viscoelastic suspensions at high volume fraction, however, it is necessary to specify how particles hydrodynamically interact when they are in close proximity, i.e., proposing a model for interparticle lubrication as done for Newtonian fluids in Prasanna Kumar et al. (2021) for example. Currently, there are no numerical techniques available for analysing dense viscoelastic suspensions since there is a lack of generalizations regarding non-Newtonian lubrication forces acting between particles. Novel lubrication force models would facilitate the development of precise and efficient reduced-order methods of the Lubrication Dynamics (LD) or Discrete Element Method (DEM) type (Prasanna Kumar et al., 2021; Ruiz-López et al., 2023; Rosales-Romero et al., 2024).
@@
Analytical investigations into the squeezing flow between nearly-touching spheres have provided valuable insights into lubrication interactions across different rheological regimes. In the Newtonian case, O’Neill (1969) presented the exact solutions for the asymmetrical slow viscous flows of an infinite fluid caused by either the rotation or by the translation of spheres along directions perpendicular to their line of centres, and Jeffrey (1982) employed an asymptotic analysis to evaluate the force between two spheres of unequal radii approaching each other with equal and opposite velocities under Stokes conditions. Extending this framework to non-Newtonian settings, Rodin (1996) derived asymptotic solutions for Power-Law fluids, showing that resistance is dominated by the gap flow only when the Power-Law index is $m\geq 1/3$, while for $m<1/3$, outer-region contributions become significant. Further, recent advancement has been done by employing both a bi-viscous shear thinning (Vázquez-Quesada and Ellero, 2016) and shear-thickening fluid model (Vázquez-Quesada et al., 2018) to obtain approximate expressions for the generalized lubrication force between adjacent spheres in both fluids.
@@
In contrast to inelastic lubricating fluids, studies focusing on viscoelastic media are comparatively limited. Chun-hui et al. (2004) used a perturbation method based on Reynolds’ lubrication theory to estimate forces between rigid spheres in a second-order fluid, while Dandekar and Ardekani (2021) employed a perturbation expansion in the Deborah number to derive viscoelastic contributions to hydrodynamic forces, including lift components along the line connecting sphere centres. Recent work by Zheng et al. (2023, 2024) applied a lubrication approach and order-of-magnitude analysis to squeezing flows in upper-convected Maxwell fluids, offering estimates for pressure, shear stress, and normal stress. Ruangkriengsin et al. (2024) also examined Oldroyd-B fluids and showed that viscoelastic effects can either enhance or reduce resistance depending on the order of approximation and motion direction. Despite these advances, most studies assume that the velocity field remains parabolic and quasi-steady, as in Newtonian (Phan-Thien and Tanner, 1984; Phan-Thien et al., 1985; Zheng et al., 2023) or shear-thinning fluids (McClelland and Finlayson, 1983; Bair, 2015), thereby neglecting direct modifications to flow kinematics arising from viscoelastic normal stress differences. This simplification, while practical, may limit model accuracy in regimes where elastic effects significantly influence flow structure.
@@
In this study, we present a semi-analytical solution for the lubrication force between two rigid spheres approaching each other in a viscoelastic fluid, modelled using the Oldroyd-B (Oldroyd and Wilson, 1950) constitutive equation under small Deborah numbers. Unlike prior approaches assuming Newtonian velocity fields, our formulation directly solves the fully coupled mass, momentum, and constitutive equations using lubrication theory and order-of-magnitude analysis. We consider both steady and transient regimes under a constant approach departure velocity without prescribing a velocity profile a priori. Our findings enable future numerical LD/DEM simulations of dense particle suspensions within Boger-type viscoelastic fluids. The structure of the paper is the following: in Section 2, the flow settings, the conservation and constitutive equations are described for the approaching-spheres flow system under consideration. In Section 3 the steady-state problem is discussed, whilst Section 4 focuses on the transient-state. In Sections 5 predictions with a real Boger fluid characterized by Oldroyd-B model is made. Finally, conclusions are offered in Section 6.
@@ 2.1 Squeezing flow between two spherical particles
Figure 1: Schema of two spheres approaching with relative velocity, $V$. The minimum distance between the sphere’ contours is $h_{0}$.
@@ 2.2 Conservation equations and constitutive equation
The mass and momentum conservation equations for an incompressible fluid ignoring the body force are, respectively:
$\nabla\cdot\boldsymbol{v}=0,$ (1)
$\rho\frac{D\boldsymbol{v}}{Dt}=\nabla\cdot\boldsymbol{\Pi}.$ (2)
@@
where $p$ is the isotropic pressure, $\boldsymbol{I}$ is the identity tensor, and the remaining two terms compose the extra-stress. The extra-stress can be decomposed into two parts: a purely-viscous component, $2\eta_{s}\boldsymbol{D}$, that corresponds to the Newtonian solvent contribution to the total stress ($\eta_{s}$ is the solvent viscosity and $\boldsymbol{D}$ is the rate-of-deformation tensor, defined as $\boldsymbol{D}=\frac{1}{2}\left(\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T}\right)$), and the viscoelastic-stress, $\boldsymbol{\tau}$, whose evolution is governed by the upper-convected Maxwell model, as follows:
@@
Here $\lambda_{1}$ is the relaxation time, $\eta_{p}$ is the polymeric viscosity, and $\overset{\bigtriangledown}{\boldsymbol{\tau}}$ denotes the upper-convected derivative, given by:
$\overset{\bigtriangledown}{\boldsymbol{\tau}}=\frac{\partial\boldsymbol{\tau}}{\partial t}+\boldsymbol{v}\cdot\nabla\boldsymbol{\tau}-\nabla\boldsymbol{v}^{T}\cdot\boldsymbol{\tau}-\boldsymbol{\tau}\cdot\nabla\boldsymbol{v}.$ (5)
Additional parameters are introduced for viscosity ratios: the total viscosity of the viscoelastic solution is $\eta_{0}=\eta_{s}+\eta_{p}$. From here, a solvent-fraction measure $\beta=\frac{\eta_{s}}{\eta_{0}}$ may be defined to account for the relative strength of the solute-to-solvent contributions. From here, one may recover the solvent viscosity $\eta_{s}=\eta_{0}\beta$ and the polymeric viscosity $\eta_{p}=\eta_{0}(1-\beta)$.
@@ 2.3 Problem statement
the momentum balance equations in $r-$direction becomes:
$\rho\left(\frac{\partial v_{r}}{\partial t}+v_{r}\frac{\partial v_{r}}{\partial r}+v_{z}\frac{\partial v_{r}}{\partial z}\right)=-\frac{\partial p}{\partial r}+\eta_{s}\left[\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial(rv_{r})}{\partial r}\right)+\frac{\partial^{2}v_{r}}{\partial z^{2}}\right]+\left(\frac{1}{r}\frac{\partial\left(r\tau_{rr}\right)}{\partial r}+\frac{\partial\tau_{rz}}{\partial z}-\frac{\tau_{\theta\theta}}{r}\right),$ (7)
@@
$\rho\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+v_{z}\frac{\partial v_{z}}{\partial z}\right)=-\frac{\partial p}{\partial z}+\eta_{s}\left[\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial v_{z}}{\partial r}\right)+\frac{\partial^{2}v_{z}}{\partial z^{2}}\right]+\left(\frac{1}{r}\frac{\partial\left(r\tau_{rz}\right)}{\partial r}+\frac{\partial\tau_{zz}}{\partial z}\right).$ (8)
Because of the symmetry in the $\theta-$direction, the associated stress equations are inherently satisfied. To close the problem, we impose no-slip boundary conditions over the surface of the spheres, as follows:

### v2

@@ Abstract
Viscoelastic fluids play a critical role in various engineering and biological applications, where their lubrication properties are strongly influenced by relaxation times ranging from microseconds to minutes. Although the lubrication mechanism for Newtonian fluids is well-established, its extension into viscoelastic materials—particularly under squeezing flow conditions—remains less explored. This study presents a semi-analytical solution for the lubrication force between two spheres approaching in a Boger fluid under small Deborah numbers. Unlike previous works that assumed a Newtonian velocity field, we derive the velocity profile directly from the mass-momentum conservation and Oldroyd-B constitutive equations using lubrication theory and order-of-magnitude analysis techniques. Under steady-state conditions, viscoelasticity induces a marginal increase in the surface-to-surface normal force as a result of the increased pressure required to overcome the resistance originating from the first normal-stress difference. Transient analyses reveal that the normal lubrication force is bounded by two Newtonian plateaus and is non-symmetric as the spheres approach or separate. Our findings highlight the role of viscoelasticity in improving load capacity and provide new insights for modelling dense particle suspensions in Boger fluids, where short-range interactions dominate.
@@ 1 Introduction
Viscoelastic fluids are fundamental in various technological and biological contexts involving close-contact flows, such as lubrication of sliding surfaces, synthetic transport in living organisms, and synovial joint dynamics (Ahmed and Biancofiore, 2021; Urry, 1993; Ateshian, 2009). Their defining feature—a wide range of relaxation times—leads to complex non-linear responses governed by normal-stress differences (Zhang and Li, 2005; Nam et al., 2015). Despite extensive studies on Newtonian suspensions (Hoffman, 1972; Zarraga et al., 2000; Scirocco et al., 2020; Guazzelli and Pouliquen, 2018), suspensions in viscoelastic media remain less explored. Experiments with viscoelastic fluids, including Boger fluids characterised by a nearly constant shear viscosity and pronounced elasticity (Oldroyd, 1950; Boger, 1977), reveal distinctive behaviours, such as early shear thickening, complex stress scaling, and matrix–microstructure interactions (Zarraga et al., 2001; Scirocco et al., 2004; Dai and Tanner, 2020). Suspensions in viscoelastic fluids also exhibit particle migration, chaining, and flow-induced structuring (Michele et al., 1997; Pasquino et al., 2013; Del Giudice et al., 2013). Microscopy studies in shear-thinning media show string-like particle alignment at low Weissenberg numbers, governed by fluid rheology and particle size (Pasquino et al., 2010). Settling experiments indicate that viscoelasticity and shear-thinning enhance migration, unlike Newtonian fluids, for which such effects are suppressed (Scirocco et al., 2020; D’Avino et al., 2012). In confined microfluidic environments, particles can self-organise into ordered structures due to coupled inertial and viscoelastic forces, underscoring the significance of confinement and non-Newtonian flow features in particle-laden systems (Villone et al., 2011; Del Giudice and D’Avino, 2024).
@@
Numerical studies on viscoelastic suspensions have largely focused on dilute to semi-dilute regimes (volume fractions up to $\phi=0.3$), offering valuable insights into particle-laden viscoelastic flows through a variety of computational frameworks. Vázquez-Quesada et al. (2019) employed Smoothed Particle Hydrodynamics (SPH), incorporating a GENERIC-based viscoelastic model, to simulate rigid particle suspensions. They observed a plateau in relative viscosity at low Deborah numbers (De), followed by shear thickening at higher De due to extensional features in the matrix. Jain and Shaqfeh (2021) used body-fitted (BF) and immersed-boundary (IB) methods to study Boger fluid suspensions, finding increases in per-particle viscosity and the first normal-stress coefficient with strain, particle volume fraction, and Weissenberg number. Interestingly, fluid stress remained independent of $\phi$, while the stresslet rose due to enhanced hydrodynamic interactions. Villone et al. (2021) investigated immersed-boundary-method numerical simulations of small-amplitude oscillatory shear in shear-thinning and viscoelastic matrices, showing inertia-induced elasticity and Carreau number–dependent shear-thinning effects. Viscoelastic moduli varied with De and viscosity ratio, but not with the Giesekus mobility factor. Several methodologies have expanded into diverse modelling tool sets: Kuron et al. (2021) proposed a lattice Boltzmann method (LBM) for Oldroyd-B fluids, preserving stress and energy in complex flows. Matsuoka et al. (2021) and Yamamoto et al. (2021) applied the smoothed profile method (SPM), effectively predicting shear thickening and accommodating complex particulate systems. Zhang and Shaqfeh (2023) combined 3D finite-volume simulations with experiments to study shear-thinning suspensions, linking reduced viscosity to diminished stresslet contributions and particle-induced stresses. Recent extensions of these frameworks include the lattice Boltzmann–smoothed profile hybrid method (LBM-SPM) by Lee (2023), and the thermally augmented SPM by Nakayama et al. (2024), which enable accurate micro-rheological analysis in thermally fluctuating viscoelastic fluids. Previous simulations of passive microrheology for a colloidal particle immersed in a thermodynamically-consistent fluctuating Oldroyd-B fluid were already presented earlier by Vázquez-Quesada et al. (2012).
@@
Despite the significant amount of numerical work performed in the dilute and semi-dilute regimes, examining dense suspensions where $\phi>0.3$ has been much less investigated. For suspensions with Newtonian matrices, recent advances in simulations of the rheology of concentrated suspensions—particularly dense, non-colloidal systems—highlight the crucial role of microstructure, contact mechanics, and interparticle forces in governing non-Newtonian flow phenomena (see Morris, 2009; Seto et al., 2013; Wyart and Cates, 2014; Tanner, 2018). To simulate suspensions with viscoelastic matrices at high volume fraction, however, it is necessary to specify how particles hydrodynamically interact when they are in close proximity, i.e., proposing a model for interparticle lubrication as done for Newtonian fluids in Prasanna Kumar et al. (2021), for example. Currently, there are no numerical techniques available for analysing dense viscoelastic suspensions since there is a lack of generalisations regarding non-Newtonian lubrication forces acting between particles. Novel lubrication force models would facilitate the development of precise and efficient reduced-order methods of the Lubrication Dynamics (LD) or Discrete Element Method (DEM) type (Prasanna Kumar et al., 2021; Ruiz-López et al., 2023; Rosales-Romero et al., 2024).
@@
Analytical investigations into the squeezing flow between nearly-touching spheres have provided valuable insights into lubrication interactions across different rheological regimes. In the Newtonian case, O’Neill (1969) presented the exact solutions for the asymmetrical slow viscous flows of an infinite fluid caused by either the rotation or by the translation of spheres along directions perpendicular to their line of centres, and Jeffrey (1982) employed an asymptotic analysis to evaluate the force between two spheres of unequal radii approaching each other with equal and opposite velocities under Stokes conditions. Extending this framework to non-Newtonian settings, Rodin (1996) derived asymptotic solutions for Power-Law fluids, showing that resistance is dominated by the gap flow only when the Power-Law index is $m\geq 1/3$, while for $m<1/3$, outer-region contributions become significant. Further recent advancement has been achieved by employing both a bi-viscous shear thinning in Vázquez-Quesada and Ellero (2016), and a shear-thickening fluid model in Vázquez-Quesada et al. (2018), to obtain approximate expressions for the generalised lubrication force between adjacent spheres in both kinds of fluids.
@@
In contrast to inelastic lubricating fluids, studies focusing on viscoelastic media are comparatively limited. Chun-Hui et al. (2004) used a perturbation method based on the Reynolds’ lubrication theory to estimate forces between rigid spheres in a second-order fluid, while Dandekar and Ardekani (2021) employed a perturbation expansion in the Deborah number to derive viscoelastic contributions to hydrodynamic forces, including lift components along the line connecting sphere centres. Recent work by Zheng et al. (2023, 2024) applied a lubrication approach and an order-of-magnitude analysis to squeezing flows in upper-convected Maxwell fluids, offering estimates for pressure, shear and normal stresses. Ruangkriengsin et al. (2024) also examined Oldroyd-B fluids and showed that viscoelastic effects can either enhance or reduce resistance depending on the order of approximation and motion direction. Despite these advances, most studies assume that the velocity field remains parabolic and quasi-steady, as in Newtonian (Phan-Thien and Tanner, 1984; Phan-Thien et al., 1985; Zheng et al., 2023) or shear-thinning fluids (McClelland and Finlayson, 1983; Bair, 2015), thereby neglecting direct modifications to flow kinematics arising from viscoelastic normal-stress differences. This simplification, while practical, may limit model accuracy in regimes where elastic effects significantly influence flow structure.
@@
In this study, we present a semi-analytical solution for the normal lubrication force between two rigid spheres approaching each other in a viscoelastic fluid, modelled using the Oldroyd-B (Oldroyd, 1950) constitutive equation under small Deborah numbers. Unlike previous approaches, which assume Newtonian velocity fields, our formulation directly solves the fully-coupled mass, momentum, and constitutive equations using lubrication theory and order-of-magnitude analysis. We consider both steady and transient regimes under constant approach or departure velocities without prescribing a velocity profile a priori. Our findings enable future numerical LD/DEM simulations of dense particle suspensions within Boger-type viscoelastic fluids. The structure of this paper is the following: in Section 2, the flow settings, and the conservation and constitutive equations are described for the approaching-spheres flow system under consideration. In Section 3, the steady-state problem is discussed, whilst Section 4 focuses on the transient state. In Section 5, predictions with a real Boger fluid characterised with the Oldroyd-B model are reported. Finally, conclusions are offered in Section 6.
@@ 2.1 Squeezing flow between two spherical particles
Figure 1: Schematics of two spheres approaching with relative velocity $V$. The minimum distance between the sphere contours is $h_{0}$.
@@ 2.2 Conservation equations and constitutive equation
The mass and momentum conservation equations for an incompressible fluid ignoring the body force are:
$\boldsymbol{\nabla}\cdot\boldsymbol{v}=0,$ (1)
$\rho\frac{D\boldsymbol{v}}{Dt}=\boldsymbol{\nabla}\cdot\boldsymbol{\Pi}.$ (2)
@@
where $p$ is the isotropic pressure, $\boldsymbol{I}$ is the identity tensor, and the remaining two terms compose the extra-stress. The extra-stress can be decomposed into two parts: a purely-viscous component, $2\eta_{s}\boldsymbol{D}$, that corresponds to the Newtonian solvent contribution to the total stress ($\eta_{s}$ is the solvent viscosity and $\boldsymbol{D}$ is the rate-of-deformation tensor, defined as $\boldsymbol{D}=\frac{1}{2}\left(\nabla\boldsymbol{v}+(\nabla\boldsymbol{v})^{T}\right)$), and the viscoelastic-stress $\boldsymbol{\tau}$ whose evolution is governed by the upper-convected Maxwell model, as follows:
@@
Here, $\lambda_{1}$ is the relaxation time, $\eta_{p}$ is the polymeric viscosity, and $\overset{\bigtriangledown}{\boldsymbol{\tau}}$ denotes the upper-convected time derivative, given by:
$\overset{\bigtriangledown}{\boldsymbol{\tau}}=\frac{\partial\boldsymbol{\tau}}{\partial t}+\boldsymbol{v}\cdot\boldsymbol{\nabla}\boldsymbol{\tau}-\boldsymbol{\nabla}\boldsymbol{v}^{T}\cdot\boldsymbol{\tau}-\boldsymbol{\tau}\cdot\boldsymbol{\nabla}\boldsymbol{v}.$ (5)
Additional parameters are introduced for viscosity ratios: the total viscosity of the viscoelastic solution is $\eta_{0}=\eta_{s}+\eta_{p}$. A solvent-fraction measure $\beta=\frac{\eta_{s}}{\eta_{0}}$ may be defined to account for the relative strength of the solute-to-solvent contributions. From here, one may recover the solvent viscosity $\eta_{s}=\eta_{0}\beta$ and the polymeric viscosity $\eta_{p}=\eta_{0}(1-\beta)$.
@@ 2.3 Problem statement
the momentum balance equation in $r-$direction becomes:
$\rho\left(\frac{\partial v_{r}}{\partial t}+v_{r}\frac{\partial v_{r}}{\partial r}+v_{z}\frac{\partial v_{r}}{\partial z}\right)=-\frac{\partial p}{\partial r}+\eta_{s}\left[\frac{\partial}{\partial r}\left(\frac{1}{r}\frac{\partial(rv_{r})}{\partial r}\right)+\frac{\partial^{2}v_{r}}{\partial z^{2}}\right]+\\ \left(\frac{1}{r}\frac{\partial\left(r\tau_{rr}\right)}{\partial r}+\frac{\partial\tau_{rz}}{\partial z}-\frac{\tau_{\theta\theta}}{r}\right),$ (7)
@@
$\rho\left(\frac{\partial v_{z}}{\partial t}+v_{r}\frac{\partial v_{z}}{\partial r}+v_{z}\frac{\partial v_{z}}{\partial z}\right)=-\frac{\partial p}{\partial z}+\eta_{s}\left[\frac{1}{r}\frac{\partial}{\partial r}\left(r\frac{\partial v_{z}}{\partial r}\right)+\frac{\partial^{2}v_{z}}{\partial z^{2}}\right]+\\ \left(\frac{1}{r}\frac{\partial\left(r\tau_{rz}\right)}{\partial r}+\frac{\partial\tau_{zz}}{\partial z}\right).$ (8)
Given the symmetry condition in the $\theta-$direction, the associated stress equations are inherently satisfied. To close the problem, we impose no-slip boundary conditions over the surface of the spheres, as follows: